Hello. ...Pete's moms age is 8 years less than 4 times Pete's age the sum of there ages are 65 what are there ages?
Answer by Maths68(1474)

Let
Present age of Mom =m
Present age of Pete =p
The sum of there ages is 65
m+p=65
m=65-p........(1)
Pete's moms age is 8 years less than 4 times Pete's age
m=4p-8........(2)
Put the value of m from (1) to (2)
65-p=4p-8
-p-4p=-8-65
-5p=-73
-5p/-5=-73/-5
p=14.6
Put the value of p in (1)
m=65-p........(1)
m=65-14.6
m=50.4
Present age of Mom =m = 50.4 years old or 50 years 4 months 24 days approximately
Present age of Pete =p = 14.6 years old or 14 years 7 months 6 days approximately